OpenStudy (anonymous):

what is the sum of a 14-term arithmetic sequence where the last term is 30 and the common difference is -5?

OpenStudy (anonymous):

do you want me to like show my work? or do you just want the answer?

OpenStudy (anonymous):

work please!

OpenStudy (anonymous):

ok so the arithmetic sequence equation is: a_n=a_1 + (n-1)d

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

since you dont have the first term to plug into the equation you plug what you have in and solve for a_1 first what do you get?

OpenStudy (anonymous):

after you get the a_1 then you can use the arithmetic Series equation: s_n =n/2[2a_1+(n-1)d] to find the sum

OpenStudy (anonymous):

so a_n=30 n=14 d=-5

OpenStudy (anonymous):

30=a_1+ (14-1)-5 30=a_1+(13)-5 30=a_1+(-65) 30=a_1-65 95=a_1

OpenStudy (anonymous):

So since you have a_1 now you plug it into the arithmetic series equation to find the sum

OpenStudy (anonymous):

so S_n=14/2[2(95)+(14+1)-5] S_n=7[190+(15)-5] S_n=190-75 S_n=115

OpenStudy (anonymous):

So the answer is 115

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

no still confused the answers is either: 875 812.5 455 422.5

OpenStudy (anonymous):

Is the problem right that you typed? cause i did my math right so is anything different?

OpenStudy (anonymous):

if the common difference is not negative it will be very different

OpenStudy (anonymous):

the question is written right, i'm confused because i got the same answer you did. but 115 isn't one of the answers i can choose from ?

OpenStudy (anonymous):

Hmm. thats odd :l none of the answers can be "none of the above"?

OpenStudy (anonymous):

no those are the only ones i can choose.

OpenStudy (anonymous):

ok hold on a second ill do my math again 805 an answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 875 if you use the other equation for the sum : S_n=n/2(a_1+a_n) S_n=14/2(95+30) S_n=7(125) S_n=875

OpenStudy (anonymous):

okay, thank you so much!

OpenStudy (anonymous):

No problem! lol